# Markov chain Monte Carlo without evaluating the target: an auxiliary variable approach

### v2

@@ Abstract
In sampling tasks, it is common for target distributions to be known up to a normalising constant. However, in many situations, evaluating even the unnormalised distribution can be costly or infeasible. This issue arises in scenarios such as sampling from the Bayesian posterior for tall datasets and the ‘doubly-intractable’ distributions. In this paper, we begin by observing that seemingly different Markov chain Monte Carlo (MCMC) algorithms, such as the exchange algorithm (Murray et al.,, 2006), PoissonMH (Zhang and De Sa,, 2019), and TunaMH (Zhang et al.,, 2020), can be unified under a simple common procedure. We then extend this procedure into a novel framework that allows the use of auxiliary variables in both the proposal and acceptance–rejection steps. We develop the theory of the new framework, applying it to existing algorithms to simplify and extend their results. Several new algorithms emerge from this framework, with improved performance demonstrated on both synthetic and real datasets.
@@ 1 Introduction
The general workflow of Markov chain Monte Carlo algorithms is similar, yet computational challenges arise in their own way. Interestingly, solutions to these unique challenges sometimes also share a common underlying principle, which often go unnoticed even by their proposers. This paper aims to explore this phenomenon by providing a unified framework for various MCMC algorithms across Bayesian inference applications. New algorithms also emerge from this framework.
We briefly recall the workflow of Metropolis–Hastings type algorithms (Metropolis et al.,, 1953; Hastings,, 1970), one of the most general and popular MCMC methods. Given a target distribution, these methods typically iterate over a two-step process: they first propose a new state from a proposal distribution, given the current state. Then, they apply an acceptance–rejection mechanism to decide if the transition to the new state should occur. While the choices of the proposal vary by algorithm, their implementation often requires evaluating the unnormalised distribution—that is, the target distribution up to a scaling factor. Several algorithms such as the Metropolis–adjusted Langevin algorithm (MALA) (Besag,, 1994) and Hamiltonian Monte Carlo (HMC) (Duane et al.,, 1987) also require evaluating the derivative of the unnormalised distribution.
In Bayesian inference, users focus on quantities associated with the posterior distribution $\pi(\theta\mid x)$, with $x$ as observed data and $\theta$ as the parameter of interest. Given a prior $\pi(\theta)$ and likelihood $p_{\theta}(x)$, the posterior distribution is $\pi(\theta\mid x)=\pi(\theta)p_{\theta}(x)/\int p_{\theta}(x)\pi(\mathrm{d}\theta)$. To ‘understand’ the typically complex posterior, MCMC is often the go-to strategy. Establishing a prior, selecting an MCMC algorithm aimed at the posterior, and examining posterior samples constitute the process of the modern Bayesian inference, with MCMC acting as the backbone. For a detailed historical background, see Martin et al., (2024).
Although MCMC needs only the unnormalised posterior, which is $\pi(\theta)p_{\theta}(x)$ as a function of $\theta$ for fixed $x$, challenges still arise when evaluating this function is expensive or not feasible. The difficulty can stem from both the $\theta$ and the $x$ side. Consider the following scenarios:
@@
Imagine that users have gathered a large dataset $x=(x_{1},\ldots,x_{N})$, where $N$ can be in the millions or billions. Each data point is independently and identically distributed following the distribution $\mathsf{p}_{\theta}$. In this case, the joint model likelihood is $p_{\theta}(x)=\prod_{i=1}^{N}\mathsf{p}_{\theta}(x_{i})$. However, executing standard Metropolis–Hastings algorithms becomes expensive, as the per-step acceptance–rejection computation requires evaluations across the full dataset. This scenario is often referred to as “tall data” (Bardenet et al.,, 2017), which represents the kind of datasets commonly found in machine learning tasks (Welling and Teh,, 2011; Ahn,, 2015).
Various approaches have been proposed to address these challenges. Here, we review three algorithms: the exchange algorithm (Algorithm 1, Murray et al., (2006)) for doubly-intractable distributions; PoissonMH (Algorithm 2, Zhang and De Sa, (2019)) and TunaMH (Algorithm 3, Zhang et al., (2020)) for tall datasets. These algorithms have demonstrated promising empirical performance compared to other existing solutions, and are most relevant to our proposed framework and new algorithms. Additional relevant studies will be discussed in Section 1.2. For completeness, the beginning of each algorithm’s pseudocode contains a section titled “promise,” which describes the technical assumptions and practical prerequisites that users need to know before executing the algorithm. It is worth noting that the posterior with $N$ i.i.d. data points satisfies this promise of PoissonMH (and TunaMH) by setting $\phi_{i}(\theta,x):=N^{-1}\log\pi(\theta)+\log\mathsf{p}_{\theta}(x_{i})$ (and $U_{i}(\theta,x):=-N^{-1}\log\pi(\theta)-\log\mathsf{p}_{\theta}(x_{i}))$.)
@@
2:  Initialise: Initial state $\theta_{0}$, proposal $q$; number of iterations $T$
@@
2:  Initialise: Initial state $\theta_{0}$; proposal $q$; hyperparameter $\lambda$; set $L:=\sum_{i=1}^{N}M_{i}$; number of iterations $T$
@@
2:  Initialise: Initial state $\theta_{0}$; proposal $q$; hyperparameter $\chi$; set $C=\sum_{i=1}^{N}c_{i}$; number of iterations $T$
@@
The similarities and differences among these algorithms are both clear. These algorithms share the goal of sampling from the posterior without directly computing the target distribution, preserving the correct stationary distribution. However, their methodologies differ. The exchange algorithm targets doubly-intractable distributions, assuming access to a simulator for synthetic data. PoissonMH and TunaMH address the ‘tall data’ problem by avoiding full dataset evaluations and impose different assumptions on the model likelihood. This prompts the question: Can these algorithms be unified under a common framework?
@@ 1.1 Our contribution
Our first contribution answers the above question affirmatively. We begin with a key observation: the exchange, PoissonMH, and TunaMH algorithms can be unified into a simple common procedure, detailed in Section 2. This procedure reveals their inherent similarities.
Furthermore, we take a step further to develop new algorithms. We notice that existing algorithms without evaluating the full posterior predominantly support random-walk-type proposals—which mix slowly in high-dimensional sampling problems. Motivated by this, our major contribution is a novel framework in Section 3.2. Thanks to this framework, we develop new gradient-based minibatch algorithms that allow users to choose a high-efficiency proposal. We develop new variants of ‘locally balanced proposals’ (Zanella,, 2020; Livingstone and Zanella,, 2022) and Stochastic Gradient Langevin Dynamics (SGLD) (Welling and Teh,, 2011) to improve the PoissonMH and TunaMH. These algorithms support gradient-based proposals while maintaining correct stationary distribution with minibatch data. They show significant efficiency improvements on synthetic and real datasets compared to their predecessors and standard algorithms like random-walk Metropolis, MALA, and HMC.
The framework itself may be of independent interest. Methodologically, it appears to be flexible and practical. It uses auxiliary variables in both the proposal and acceptance-rejection steps, encompassing previous frameworks Titsias and Papaspiliopoulos, (2018) and Section 2, which both focus on only one aspect. This offers greater potential for further algorithm design. Theoretically, the framework is linked to the theory of Markov chain comparison, an active area with recent progress (Andrieu et al.,, 2022; Power et al.,, 2024; Qin et al.,, 2023; Ascolani et al.,, 2024). Both classical (Diaconis and Saloff-Coste,, 1993; Dyer et al.,, 2006) and recent techniques are well-suited for analysing the framework or its applications. Our theoretical analysis in Section 4 can recover, strengthen, and extend several existing results in a simpler way.
@@ 1.2 Related works
For tall datasets, many minibatch MCMC algorithms reduce iteration costs. These are classified into exact and inexact methods, based on whether they maintain the target distribution as stationary. Inexact methods, such as those in Welling and Teh, (2011); Chen et al., (2014); Korattikara et al., (2014); Quiroz et al., (2018); Dang et al., (2019); Wu et al., (2022), introduce a systematic bias to trade off for implementation efficiency. Zhang et al., (2020) demonstrate that this bias can accumulate significantly, even when the bias per-step is arbitrarily small. In contrast, exact methods such as Maclaurin and Adams, (2014); Bardenet et al., (2017); Cornish et al., (2019); Zhang and De Sa, (2019); Zhang et al., (2020) maintain the correct stationary distribution but rely on assumptions about the target distribution. This paper primarily focuses on Zhang et al., (2020) and its predecessor Zhang and De Sa, (2019), as they show improved empirical performance with relatively mild assumptions. Non-reversible exact methods by Bouchard-Côté et al., (2018); Bierkens et al., (2019) are also appealing but can be challenging to implement. In addition to minibatch MCMC algorithms, divide-and-conquer algorithms by Huang and Gelman, (2005); Wang and Dunson, (2013); Neiswanger et al., (2014) are suitable for parallel implementation.
@@ 1.4 Organization
We present a common procedure that unifies the exchange algorithm, PoissonMH, and TunaMH in Section 2. Our main framework extends this further and is presented as a meta-algorithm (Algorithm 4) in Section 3. Section 3.3 introduces new algorithms as special cases of our meta-algorithm. Section 4 provides a theory for the mixing time of Algorithm 4, with applications in Section 4.4. Section 5 demonstrates the effectiveness of our new algorithms through three experiments. Section 6 concludes this paper with future directions. Additional proofs, experiment details, and extra experiments are provided in the supplementary.
@@ 2 A common procedure of existing algorithms
We provide two remarks here. An additioanl comment is provided in supplementary C. First, the novelty of our finding in this section lies in discovering the inherent connections between algorithms designed for doubly-intractable distributions and those for tall datasets. The common procedure itself is not new. In Section 2.1 of Andrieu et al., 2018a , the authors present a slightly broader framework in which they substitute $P_{\theta^{\prime}\rightarrow\theta}(\omega)$ with $P_{\theta^{\prime}\rightarrow\theta}(\varphi(\omega))$, where $\varphi$ represents any measurable involution. Therefore our framework can be viewed as taking $\varphi=\mathsf{id}$, the identity map. On the other hand, the primary goal in Andrieu et al., 2018a is to improve the quality of the estimator within the exchange algorithm while preserving the stationary distribution. Consequently, they do not consider tall datasets.
Second, it is interesting to observe that Algorithm 1, 2, and 3 use three different ways of generating the auxiliary variable. In the exchange algorithm, $\omega$ depends on $\theta^{\prime}$; in PoissonMH, it depends on $\theta$; and in TunaMH, it depends on both. While this may appear as a simple finding, it will significantly impact the design of new algorithms in Section 3.
The current common procedure already provides a convenient basis for analysing the theory behind these algorithms. Nevertheless, we choose to postpone the theoretical analysis after Section 3, where a more comprehensive framework will be introduced. Analysing the general framework allows us to simplify and improve existing results as corollaries.
@@ 3.1 Motivation: gradient-based proposal design
One simple idea is to choose $q$ as a gradient-based proposal in Section 2. However, upon second thought, one will quickly realise the challenges in implementing such a strategy: For the doubly-intractable distribution, the gradient of log-posterior involves the intractable $\nabla_{\theta}\log(Z(\theta))$. For tall datasets, computing the gradient requires evaluating the entire dataset, conflicting with the goal of reducing iteration costs in minibatch MCMC algorithms like PoissonMH and TunaMH.
Nevertheless, designing an ‘informed proposal’ remains a fascinating idea. Motivated by this, we introduce a novel framework in Section 3.2, extending Section 2. This framework leads to new algorithms using cheap proxies for the gradient to guide proposed moves.
@@ 3.2 General methodology
We present our method using notations and assumptions in algorithmic and probabilistic terms. We assume that users can access two simulators for generating auxiliary variables, ${\mathcal{S}}_{1}$ and ${\mathcal{S}}_{2}$. ${\mathcal{S}}_{1}$ takes $\theta$ and produces a random variable $\omega_{1}$. ${\mathcal{S}}_{2}$ takes $(\theta,\theta^{\prime},\omega_{1})$ and outputs another random variable $\omega_{2}$. Formally, consider two measurable spaces $(\Omega_{1},{\mathcal{F}}_{1})$ and $(\Omega_{2},{\mathcal{F}}_{2})$ with base measures $\lambda_{1}$ and $\lambda_{2}$. We define two families of probability measures: $\mathbb{P}_{\theta}$ on $\Omega_{1}$ parameterised by $\theta\in\Theta$, and $\mathbb{P}_{\theta,\theta^{\prime}}(\cdot\mid\omega_{1})$ on $\Omega_{2}$ parameterised by $(\theta,\theta^{\prime})\in\Theta\times\Theta$ and $\omega_{1}\in\Omega_{1}$. The second family represents the conditional distribution of $\omega_{2}$ given $\omega_{1}$ with parameters $(\theta,\theta^{\prime})$. For any $(\theta,\theta^{\prime})$, this defines a joint distribution on $\Omega_{1}\times\Omega_{2}$ as $\mathbb{P}_{\theta,\theta^{\prime}}(\omega_{1},\omega_{2})=\mathbb{P}_{\theta}(\omega_{1})\mathbb{P}_{\theta,\theta^{\prime}}(\omega_{2}\mid\omega_{1})$. Our assumption implies users can simulate from both the marginal distribution of $\omega_{1}$ and the conditional distribution of $\omega_{2}$. Additionally, we introduce the single-point probability space $\mathsf{NULL}:=\{\mathsf{null}\}$. Setting $\Omega_{1}$ or $\Omega_{2}$ to $\mathsf{NULL}$ indicates no corresponding auxiliary variable is generated. Lastly, we introduce a family of proposal kernels $\{Q_{\omega_{1}}\}_{\omega_{1}\in\Omega_{1}}$, where each $\omega_{1}$ is associated with a Markov transition kernel $Q_{\omega_{1}}$ on $\Theta$. We assume $Q_{\omega_{1}}(\theta,\cdot)$ has a density $q_{\omega_{1}}(\theta,\cdot)$ with respect to a base measure $\lambda$.
@@
Our intuition for introducing $\omega_{1}$ and $\omega_{2}$ is actually simple. The first auxiliary variable $\omega_{1}$ is used to determine the proposal, with its distribution depending solely on the current state $\theta$. The sampled $\omega_{1}$ and the current state $\theta$ then determine the proposed state $\theta^{\prime}$. Then, $\omega_{2}$, with distribution depending on $\theta$, $\theta^{\prime}$, and $\omega_{1}$, is generated to estimate the target ratio. For instance, $\omega_{1}$ might be a minibatch uniformly selected from a large dataset, used to estimate the gradient and influence the proposal. An example of $\omega_{2}$ could be synthetic data generated at each iteration of the exchange algorithm. Further examples will be presented shortly.
@@
The procedure of our auxiliary-variable-based meta-algorithm is detailed in Algorithm 4. For an additional illustration, refer to Figure 1. Algorithm 4 is presented in its general form, assuming the target is a generic distribution $\Pi$ on $\Theta$. In all applications discussed in this paper, $\Pi(\cdot)$ is taken as the posterior distribution $\pi(\cdot\mid x)$. It is important to note that, evaluating the ratio is in Step 6 can be much cheaper than evaluating $\Pi(\theta^{\prime})/\Pi(\theta_{t})$. The costly part of $\Pi(\theta^{\prime})/\Pi(\theta_{t})$ is often offset by the ratio $\mathbb{P}_{\theta^{\prime},\theta_{t}}(\omega_{1},\omega_{2})/\mathbb{P}_{\theta_{t},\theta^{\prime}}(\omega_{1},\omega_{2})$ by design.
@@
Let $\mathbb{P}_{\mathsf{aux}}(\cdot,\cdot)$ be the transition kernel of the Markov chain defined in Algorithm 4. We have $\mathbb{P}_{\mathsf{aux}}(\cdot,\cdot)$ is reversible respect to $\Pi$.  [^1] The ‘acceptance–rejectionion’ mechanism ($\min\{1,r\}$ in Step 7) can be generalised to accept $\theta^{\prime}$ with a probability of $a(r)$, where $a:[0,\infty)\rightarrow[0,1]$ is any function satisfying $a(t)=ta(1/t)$. The reversibility claimed here still holds. See supplementary A.12 for the proof.
@@
1:  Initialise: Initial state $\theta_{0}$, auxiliary-based proposal $\{q_{\omega_{1}}\}_{\omega_{1}\in\Omega_{1}}$; number of iterations $T$; target distribution $\Pi(\theta)$
@@
From now on, we will always assume $\Pi$ is the posterior $\pi(\cdot\mid x)$ without further specification. When both $\Omega_{1}=\Omega_{2}=\mathsf{NULL}$, then no auxiliary variable is generated. Algorithm 4 reduces to the standard Metropolis–Hastings algorithm. We will now briefly discuss a few other possibilities of choosing $\omega_{1}$ and $\omega_{2}$.
@@
If $\Omega_{1}=\mathsf{NULL}$, Algorithm 4 does not utilise any auxiliary variable for the proposal distribution. The acceptance ratio in Step 6 of Algorithm 4 simplifies to
@@
This scenario recovers the previous common procedure discussed in Section 2, which in turn includes the exchange algorithm, PoissonMH, and TunaMH (Algorithm 1, 2, 3).
@@
In this case, $\omega_{1}$ can assist in designing the proposal distribution. This approach can sometimes also help estimate the target ratio and reduce per-iteration costs through careful design, as discussed in Section 3.3.1. This includes the auxiliary Metropolis-Hastings sampler (Section 2.1 of Titsias and Papaspiliopoulos, (2018)) as a special case. However, a potential limitation is that $\omega_{1}$ is generated based solely on the current state $\theta_{t}$. Incorporating information from the proposed state $\theta^{\prime}$ could further improve the target ratio estimator.

### v3

@@ Abstract
In sampling tasks, it is common for target distributions to be known up to a normalizing constant. However, in many situations, even evaluating the unnormalized distribution can be costly or infeasible. This issue arises in scenarios such as sampling from the Bayesian posterior for tall datasets and the ‘doubly-intractable’ distributions. In this paper, we begin by observing that seemingly different Markov chain Monte Carlo (MCMC) algorithms, such as the exchange algorithm, PoissonMH, and TunaMH, can be unified under a simple common procedure. We then extend this procedure into a novel framework that allows the use of auxiliary variables in both the proposal and the acceptance–rejection step. Several new MCMC algorithms emerge from this framework that utilize estimated gradients to guide the proposal moves. They have demonstrated significantly better performance than existing methods on both synthetic and real datasets. Additionally, we develop the theory of the new framework and apply it to existing algorithms to simplify and extend their results.
@@ 1 Introduction
The general workflow of Markov chain Monte Carlo algorithms is similar, yet computational challenges arise in their own way. Interestingly, solutions to these unique challenges sometimes also share a common underlying principle, which often go unnoticed even by their proposers. This paper presents an auxiliary variable-based framework that integrates several MCMC algorithms in Bayesian inference applications. This unified perspective also inspires the development of novel algorithms with improved performance.
We briefly recall the workflow of Metropolis–Hastings type algorithms (Metropolis et al.,, 1953; Hastings,, 1970), one of the most general and popular MCMC methods. Given a target distribution, these methods typically iterate over a two-step process: they first propose a new state from a proposal distribution given the current state. Then, they apply an acceptance–rejection mechanism to decide if the transition to the new state should occur. Although the choices of the proposal vary by algorithm, their implementation often requires evaluating the unnormalized distribution, that is, the target distribution up to a scaling factor. Several algorithms such as the Metropolis–adjusted Langevin algorithm (MALA) (Besag,, 1994) and Hamiltonian Monte Carlo (HMC) (Duane et al.,, 1987) also require evaluating the derivative of the unnormalized distribution.
In Bayesian inference, users focus on quantities associated with the posterior distribution $\pi(\theta\mid x)$, with $x$ as observed data and $\theta$ as the parameter of interest. Given a prior $\pi(\theta)$ and likelihood $p_{\theta}(x)$, the posterior distribution is $\pi(\theta\mid x)=\pi(\theta)p_{\theta}(x)/\int p_{\theta}(x)\pi(\mathrm{d}\theta)$. To ‘understand’ the typically complex posterior, MCMC is often the go-to strategy. For a detailed historical background, see Martin et al., (2024).
Although MCMC needs only the unnormalized posterior, which is $\pi(\theta)p_{\theta}(x)$ as a function of $\theta$ for fixed $x$, computational challenges can still arise from both the $\theta$ and the $x$ side. Consider the following scenarios:
@@
Imagine that users have gathered a large dataset $x=(x_{1},\ldots,x_{N})$, where $N$ can be in the millions or billions. Each data point is independently and identically distributed following the distribution $\mathsf{p}_{\theta}$. In this case, the joint model likelihood is $p_{\theta}(x)=\prod_{i=1}^{N}\mathsf{p}_{\theta}(x_{i})$. However, executing standard Metropolis–Hastings algorithms becomes expensive, as the per-step acceptance–rejection computation requires evaluations across the full dataset. This scenario is called “tall data” (Bardenet et al.,, 2017), which is common in machine learning (Welling and Teh,, 2011; Ahn,, 2015).
Various approaches have been proposed to address these challenges. Here, we review three algorithms: the exchange algorithm (Algorithm 1, Murray et al., (2006)) for doubly-intractable distributions; PoissonMH (Algorithm 2, Zhang and De Sa, (2019)) and TunaMH (Algorithm 3, Zhang et al., (2020)) for tall datasets. These algorithms have demonstrated promising empirical performance compared to other existing solutions, and are most relevant to our proposed framework and new algorithms. Additional relevant studies will be discussed in Section 1.2. For completeness, the beginning of each algorithm’s pseudocode contains a section titled “promise,” which describes the technical assumptions and practical prerequisites that users need to know before executing the algorithm. Two points merit attention. First, the posterior with $N$ i.i.d. data points satisfies this promise of PoissonMH (and TunaMH) by setting $\phi_{i}(\theta,x):=N^{-1}\log\pi(\theta)+\log\mathsf{p}_{\theta}(x_{i})$ (and $U_{i}(\theta,x):=-N^{-1}\log\pi(\theta)-\log\mathsf{p}_{\theta}(x_{i}))$.) Second, the naïve implementation of the Poisson minibatch sampling (Step 5 in Algorithm 2, Step 6 in Algorithm 3) has a cost linear with the total data size, since each data point is associated with a Poisson variable. However, as highlighted in Zhang and De Sa, (2019); Zhang et al., (2020), the ‘Poisson thinning’ trick can be used to implement Poisson minibatch sampling much more efficiently. Therefore, its per-step cost is much lower than both the naïve implementation and full-batch methods. See Supplementary B for implementation details and proofs.
@@
2:  Initialize: Initial state $\theta_{0}$, proposal $q$; number of iterations $T$
@@
2:  Initialize: Initial state $\theta_{0}$; proposal $q$; hyperparameter $\lambda$; set $L:=\sum_{i=1}^{N}M_{i}$; number of iterations $T$
@@
2:  Initialize: Initial state $\theta_{0}$; proposal $q$; hyperparameter $\chi$; set $C=\sum_{i=1}^{N}c_{i}$; number of iterations $T$
@@
The similarities and differences among these algorithms are clear. All these algorithms aim to sample without explicitly calculating the target distribution, while still preserving the correct stationary distribution. However, they vary significantly in their motivations, methodologies, technical assumptions, and algorithmic details. This prompts the question: Can these algorithms be unified under a common framework?
@@ 1.1 Our contribution
Our first contribution answers the above question affirmatively. We begin with a key observation: the exchange, PoissonMH, and TunaMH algorithms can be unified into a simple common procedure, detailed in Section 2.
Furthermore, we notice that existing algorithms without evaluating the full posterior predominantly support random-walk-type proposals—which mix slowly in high-dimensional sampling problems. Motivated by this, our major contribution is a novel framework in Section 3.2. Thanks to this framework, we develop new gradient-based minibatch algorithms that allow users to choose a high-efficiency proposal based on gradient information. These algorithms support gradient-based proposals while maintaining correct stationary distribution with minibatch data.
Our framework (presented in meta-algorithm 4) uses auxiliary variables in both the proposal and the acceptance-rejection step. These auxiliary variables are typically generated in a computationally efficient manner and are used to estimate costly quantities, such as required quantities associated with the proposal and the target ratio.
At a high-level, we wish to convey the information that higher-performance MCMC can be designed by replacing the expensive quantities in both the proposal and the acceptance-rejection step with cheap estimates, while still preserving the stationary distribution through appropriate design. This information is supported by our new algorithms (Algorithm 5, 6) within our framework. They have demonstrated significantly improved performance compared to their predecessors and standard MCMC methods (including random-walk Metropolis, MALA, Barker (in the sense of Livingstone and Zanella, (2022)) and HMC.)
Furthermore, our framework can help remove the systematic bias present in existing minibatch gradient-based MCMC methods that do not converge to the target distribution. An important example is the celebrated SGLD algorithm, which is known to be efficient but exhibit systematic bias, and correcting this bias using minibatch data has remained an open challenge. To quote from Welling and Teh, (2011): “ Interesting directions of future research includes deriving a MH rejection step based on mini-batch data $\ldots$ Our Tuna–SGLD (Algorithm 6) provides a solution to this problem under mild assumptions.
While the use of auxiliary variables is well-established in the MCMC algorithm design literature, our framework seems to be more comprehensive than existing approaches. It incorporates auxiliary variables in both the proposal and acceptance-rejection steps, whereas previous frameworks, such as Titsias and Papaspiliopoulos, (2018), many existing algorithms, and the methods discussed in Section 2, focus on one of these two parts. Theoretically, the framework is linked to the theory of Markov chain comparison, an active area with recent progress (Andrieu et al.,, 2022; Power et al.,, 2024; Qin et al.,, 2023; Ascolani et al.,, 2024). Our theoretical analysis in Section 4 can recover, strengthen, and extend several existing results in a simpler way.
@@ 1.2 Related works
For tall datasets, many minibatch MCMC algorithms reduce iteration costs. These are classified into exact and inexact methods, based on whether they maintain the target distribution as stationary. Inexact methods, such as those in Welling and Teh, (2011); Chen et al., (2014); Korattikara et al., (2014); Quiroz et al., (2018); Dang et al., (2019); Wu et al., (2022), introduce a systematic bias to trade off for implementation efficiency. Zhang et al., (2020) demonstrate that this bias can accumulate significantly, even when the bias per-step is arbitrarily small. In contrast, exact methods such as Maclaurin and Adams, (2014); Bardenet et al., (2017); Cornish et al., (2019); Zhang and De Sa, (2019); Zhang et al., (2020) maintain the correct stationary distribution but rely on assumptions about the target distribution. This paper primarily focuses on Zhang et al., (2020) and its predecessor Zhang and De Sa, (2019), as they show improved empirical performance with relatively mild assumptions. Non-reversible exact methods by Bouchard-Côté et al., (2018); Bierkens et al., (2019) are also appealing but can be challenging to implement. Besides minibatch MCMC algorithms, divide-and-conquer algorithms by Huang and Gelman, (2005); Wang and Dunson, (2013); Neiswanger et al., (2014) are suitable for parallel implementation.
Following the initial release of our work, Prado et al., (2024) introduced a new minibatch MCMC sampler that uses control variates to improve performance.
@@ 1.4 Organization
We present a common procedure that unifies the exchange algorithm, PoissonMH, and TunaMH in Section 2. Our main framework extends this further and is presented as meta-algorithm 4 in Section 3. Section 3.3 introduces new algorithms as special cases of our meta-algorithm. Section 4 provides a theory for the mixing time of Algorithm 4, with applications in Section 4.4. Section 5 demonstrates the effectiveness of our new algorithms through three experiments. Section 6 concludes this paper with future directions. Additional proofs, experiment details, and extra experiments are provided in the supplementary.
@@ 2 A common procedure of existing algorithms
We provide two remarks here. First, the novelty of our finding in this section lies in discovering the inherent connections between algorithms designed for doubly-intractable distributions and those for tall datasets. The common procedure itself is not new. In Section 2.1 of Andrieu et al., 2018a , the authors present a slightly broader framework in which they substitute $P_{\theta^{\prime}\rightarrow\theta}(\omega)$ with $P_{\theta^{\prime}\rightarrow\theta}(\varphi(\omega))$, where $\varphi$ represents any measurable involution. Our framework can be viewed as taking $\varphi$ to be the identity map. On the other hand, the primary goal in Andrieu et al., 2018a is to improve the quality of the estimator within the exchange algorithm. Consequently, they do not consider tall datasets.
Second, Algorithm 1, 2, and 3 use different ways of generating the auxiliary variable. In the exchange algorithm, $\omega$ depends on $\theta^{\prime}$; in PoissonMH, it depends on $\theta$; and in TunaMH, it depends on both. This observation will impact our new design in Section 3.
@@ 3.1 Motivation: gradient-based proposal design
One simple idea is to choose $q$ as a gradient-based proposal in Section 2. However, upon second thought, one will quickly realize the challenges in implementing such a strategy: For the doubly-intractable distribution, the gradient of log-posterior involves the intractable $\nabla_{\theta}\log(Z(\theta))$. For tall datasets, computing the gradient requires evaluating the entire dataset, conflicting with the goal of reducing iteration costs in minibatch MCMC algorithms like PoissonMH and TunaMH.
Nevertheless, designing an ‘gradient-based proposal’ remains a fascinating idea. Motivated by this, we introduce a novel framework in Section 3.2, extending Section 2. This framework introduces new algorithms that use computationally efficient gradient estimators to guide proposed moves.
@@ 3.2 General methodology
We present our method using notations and assumptions in algorithmic and probabilistic terms. We assume that users can access two simulators for generating auxiliary variables, ${\mathcal{S}}_{1}$ and ${\mathcal{S}}_{2}$. ${\mathcal{S}}_{1}$ takes $\theta$ and produces a random variable $\omega_{1}$. ${\mathcal{S}}_{2}$ takes $(\theta,\theta^{\prime},\omega_{1})$ and outputs another random variable $\omega_{2}$. Formally, consider two measurable spaces $(\Omega_{1},{\mathcal{F}}_{1})$ and $(\Omega_{2},{\mathcal{F}}_{2})$ with base measures $\lambda_{1}$ and $\lambda_{2}$. We define two families of probability measures: $\mathbb{P}_{\theta}$ on $\Omega_{1}$ parameterized by $\theta\in\Theta$, and $\mathbb{P}_{\theta,\theta^{\prime}}(\cdot\mid\omega_{1})$ on $\Omega_{2}$ parameterized by $(\theta,\theta^{\prime})\in\Theta\times\Theta$ and $\omega_{1}\in\Omega_{1}$. The second family represents the conditional distribution of $\omega_{2}$ given $\omega_{1}$ with parameters $(\theta,\theta^{\prime})$. For any $(\theta,\theta^{\prime})$, this defines a joint distribution on $\Omega_{1}\times\Omega_{2}$ as $\mathbb{P}_{\theta,\theta^{\prime}}(\omega_{1},\omega_{2})=\mathbb{P}_{\theta}(\omega_{1})\mathbb{P}_{\theta,\theta^{\prime}}(\omega_{2}\mid\omega_{1})$. Our assumption implies users can simulate from both the marginal distribution of $\omega_{1}$ and the conditional distribution of $\omega_{2}$. Additionally, we introduce the single-point probability space $\mathsf{NULL}:=\{\mathsf{null}\}$. Setting $\Omega_{1}$ or $\Omega_{2}$ to $\mathsf{NULL}$ indicates no corresponding auxiliary variable is generated. Lastly, we introduce a family of proposal kernels $\{Q_{\omega_{1}}\}_{\omega_{1}\in\Omega_{1}}$, where each $\omega_{1}$ is associated with a Markov transition kernel $Q_{\omega_{1}}$ on $\Theta$. We assume $Q_{\omega_{1}}(\theta,\cdot)$ has a density $q_{\omega_{1}}(\theta,\cdot)$ with respect to a base measure $\lambda$.
@@
Our intuition for introducing $\omega_{1}$ and $\omega_{2}$ is actually simple. The first auxiliary variable $\omega_{1}$ is used to determine the proposal, with its distribution depending solely on the current state $\theta$. The sampled $\omega_{1}$ and the current state $\theta$ then determine the proposed state $\theta^{\prime}$. Then, $\omega_{2}$, with distribution depending on $\theta$, $\theta^{\prime}$, and $\omega_{1}$, is generated to estimate the target ratio. For instance, $\omega_{1}$ might be a minibatch uniformly selected from a large dataset, used to estimate the gradient and influence the proposal. An example of $\omega_{2}$ could be synthetic data generated in the exchange algorithm. Further examples will be presented shortly.
@@
Our meta-algorithm is detailed in Algorithm 4. See also Figure 1 for an additional illustration. Algorithm 4 assumes the target is a generic distribution $\Pi$ on $\Theta$. In all applications discussed in this paper, $\Pi(\cdot)$ is the posterior distribution $\pi(\cdot\mid x)$. It is important to note that, evaluating the ratio in Step 6 can be much cheaper than evaluating $\Pi(\theta^{\prime})/\Pi(\theta_{t})$. The costly part of $\Pi(\theta^{\prime})/\Pi(\theta_{t})$ is often offset by $\mathbb{P}_{\theta^{\prime},\theta_{t}}(\omega_{1},\omega_{2})/\mathbb{P}_{\theta_{t},\theta^{\prime}}(\omega_{1},\omega_{2})$ by design.
@@
Let $\mathbb{P}_{\mathsf{aux}}(\cdot,\cdot)$ be the transition kernel of the Markov chain defined in Algorithm 4. We have $\mathbb{P}_{\mathsf{aux}}(\cdot,\cdot)$ is reversible respect to $\Pi$.  [^1] The ‘acceptance–rejectionion’ mechanism ($\min\{1,r\}$ in Step 7) can be generalized to accept $\theta^{\prime}$ with a probability of $a(r)$, where $a:[0,\infty)\rightarrow[0,1]$ is any function satisfying $a(t)=ta(1/t)$. The reversibility claimed here still holds. See supplementary A.12 for the proof.
@@
1:  Initialize: Initial state $\theta_{0}$, auxiliary-based proposal $\{q_{\omega_{1}}\}_{\omega_{1}\in\Omega_{1}}$; number of iterations $T$; target distribution $\Pi(\theta)$
@@
From now on, we will always assume $\Pi$ is the posterior $\pi(\cdot\mid x)$. When both $\Omega_{1}=\Omega_{2}=\mathsf{NULL}$, then no auxiliary variable is generated. Algorithm 4 reduces to the standard Metropolis–Hastings algorithm. We will now briefly discuss several other possibilities.
@@
If $\Omega_{1}=\mathsf{NULL}$, Algorithm 4 does not utilize any auxiliary variable for the proposal distribution. The acceptance ratio in Step 6 of Algorithm 4 simplifies to
@@
This recovers Section 2, which in turn includes the exchange algorithm, PoissonMH, and TunaMH (Algorithm 1, 2, 3).
@@
In this case, $\omega_{1}$ can assist in designing the proposal distribution. This approach can sometimes also help estimate the target ratio and reduce per-iteration costs through careful design, as discussed in Section 3.3.1. This includes the auxiliary Metropolis-Hastings sampler (Section 2.1 of Titsias and Papaspiliopoulos, (2018)) as a special case. However, a potential limitation is that $\omega_{1}$ only relies on the current state $\theta_{t}$. Incorporating information from the proposed state $\theta^{\prime}$ could further improve the target ratio estimator.